40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 237     Entry time: Mon Jul 26 22:14:39 2010
Author: Megan 
Type: DailyProgress 
Category:  
Subject: Daily Update 

Today I aligned the AOM so that I could align the analyzer cavity to lock the cavity. We already aligned and locked the PMC last week so we had a beam to work with. The AOM is now aligned and the beam coming from it is aligned into the analyzer cavity. We scanned the cavity to find the mode we want and currently the output of the photocell is being mixed with the local oscillator to find the error signal. It's not quite how it should be right now, so we will look at it tomorrow to figure out what is wrong. Then we can connect it to the PDH servo and lock the cavity!

I also did some cable-management. They are now neatly organized and held in place by a large number of zip ties so they don't get in the way of everything else! The cameras are now connected so that we can see the beam from the PMC, analyzer cavity, and reference cavity.

ELOG V3.1.3-